OpenStudy (anonymous):

**Calc 2 Help Needed** Solve the initial value problem below: (t^2 + 2t)dx/dt = 2x + 5 with t, x>0 and x(1) = 1

OpenStudy (bahrom7893):

(t^2 + 2t)dx/dt = 2x + 5, multiply both sides by dt: (t^2 + 2t)dx = (2x + 5)dt, divide both sides by t^2+2t dx = (2x + 5)dt/(t^2 + 2t), divide both sides by 2x+5 dx/(2x+5) = dt/(t^2+2t)

OpenStudy (bahrom7893):

Integrate both sides: Integral(dx/(2x+5)) = Integral(dt/(t^2+2t)) For the left integral use substitution u = 2x+5, du = 2xdx For the right integral use partial fractions: A/t + B/(t+2) = 1/(t*(t+2))

OpenStudy (bahrom7893):

solve for both integrals and tell me what u get. If u get lost, ask. I'll be working on my own hw meanwhile...

OpenStudy (anonymous):

YOu still there??